John and Ann are having a birthday party. John invited 14 people and will serve 3 cakes. Ann invited 28 people and is serving 7 cakes. At whose party will the guests get the most cake?
3 cakes to 14 people = 3/14 = .21 piece of cake p/person
7 cakes to 28 people = 7/28 = .25 piece of cake p/person
The most cake p/guest will be served at Ann's party.
This right?

Answer by checkley77(12844)

3 cakes to 14 people = 3/14 = .21 piece of cake p/person
7 cakes to 28 people = 7/28 = .25 piece of cake p/person
The most cake p/guest will be served at Ann's party.
CORRECT.